OpenStudy (anonymous):

Solve the equation and check your solution 2x/5 + 9 = 1/7

OpenStudy (anonymous):

solving for x?

OpenStudy (anonymous):

2x = 1/7 - 9, 2x = -62/7, x = --62/14, x = - 31/7.... booom

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

lol missed that 5 didn't you

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

2x/5 = -62/7, 14x = - 62 times 5, 14x = - 130, 7x = -65, x = - 65/7.....boom

OpenStudy (anonymous):

see yah dummes

OpenStudy (anonymous):

alrighty then...doesn't match the work I have...but okay.